OpenStudy (anonymous):

Solve by factoring the GCF: 16x^2-4x=0

OpenStudy (anonymous):

let's break both 16x^2 and 4x down 16x^2 = 2 * 2 * 2 * 2 * x * x 4x = 2 * 2 * x which are the same?

OpenStudy (anonymous):

2*2*x

OpenStudy (anonymous):

GCF is 4x 4x(4x - 1) = 0 4x = 0 or 4x = 1 can you continue?

OpenStudy (anonymous):

2 * 2 * x = 4x take 4x out~ 16x^2 - 4x = 0 4x(4x - 1) = 0 4x = 0 4x - 1 = 0 ~

OpenStudy (anonymous):

thank you!